General Meeting Summary 12/7/16

From cua_phy
Jump to navigation Jump to search

DETECTOR UPDATES

  • PbWO4 light yield (Richard/Salina/Rishabh)
  • Investigating light loss compared to earlier measurements. Currently checking
  • Pad misalignment at connection crystal/PMT (see photograph) - re-measuring light yield for J19
  • Stepper motor operation - it's not working at the moment, so all plots from this and previous meetings were really functions of time
  • thermal annealing of J21 started today
  • Also planning to check:
  • PMT gain
  • Optical coupling crystal/PMT
  • Analysis code details


  • Alternative crystal readout (Andres)
  • Got a good signal from borrowed HPS setup
  • For readout with ADC need to adjust the gate, but borrowed setup from HPS does not seem to have that feature. May also need to invert the gate for readout
  • Plan is to continue work to measure light yield with APDs and compare to PMT


  • NPS PMT shielding concept test (Arthur)
  • PMT wrapped with 500um mu-metal (Ni-Fe-Mo alloy)
  • As field is applied, signal preserved shape independent of direction of applied field
  • For transverse field lose ~20% of signal at 30 Gauss - need to check the strength of anticipated transverse fields at location of PMTs in NPS
  • For axial field lose ~20% of signal at 5 Gauss


  • Aerogel detector tray exchange (Arthur)
  • Cover for SP-11 tray transportation can be done after 10 December - Walter's group
  • Still need to install 8th layer of aerogel in SP-11 tray
  • Discussion about tray exchange before Christmas break ongoing


  • DIRC GluEx (Greg)
  • Characterization of optical properties of "cookies" ongoing - transmittance measurements with Lambda950 at CUA. Collaboration with MIT group at JLab
  • Good progress on monitoring system to transport BaBar DIRC boxes to JLab. Initial estimate is that first box will be transported in Spring 2017. Setup of the system will be done in Hall D space at JLab. Collaboration with GSI group.


  • Pixel detectors (Rachel)
  • Preparations for detector work over Christmas break ongoing - CERN will be closed 22 Dec-4 Jan
  • Integration of pixel detectors progresses well. Leak in one of the detectors was found and is currently being repaired. Cable replacements on the detectors are also currently taken care of.
  • Shipment of instrumentation from Nebraska is delayed until spring 2017


TCS PROPOSAL

  • Theory projections are essential for the proposal - need to check again with Marie and perhaps Michel
  • Acceptance studies are also needed - this part progressing well